OpenStudy (anonymous):

Find the domain of the function and express the answer in interval notation: f(x)=sqrt(4x-16)

OpenStudy (anonymous):

\[4x-16\geq 0\] \[4x\geq 16\] \[x\geq4\]

OpenStudy (anonymous):

because you cannot take the square root of a negative number

OpenStudy (anonymous):

i would take it that the answer to this question is (-\[\infty\], 16)

OpenStudy (anonymous):

\[\sqrt{4x-16}\ge 0\]

OpenStudy (anonymous):

the answer is \[[4,\infty)\]

OpenStudy (anonymous):

satellite is rght

OpenStudy (anonymous):

the expression inside the radical must not be negative. that is why you solve \[4x-16\geq0\] for x

OpenStudy (anonymous):

how would 4 be the answer, because i think i kinda figured out how to do this when its dividing.

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

this is \[f(x)=\sqrt{4x-16}\] yes?

OpenStudy (anonymous):

and you have to worry about the radical, because you cannot take the square root of a negative number.

OpenStudy (anonymous):

it is not the same as problems with the variable in the denominator. this time you worry about what is under the radical. it must not be negative. so we write \[4x-16\geq 0\] add 16 to get \[4x\geq16\] divide by 4 to get \[x\geq 4\]

OpenStudy (anonymous):

in interval notation it is \[[4,\infty)\]

OpenStudy (anonymous):

ok, i can see how you got that. i think that i just dont get how everything changes depending on how the problem is set up.